OpenStudy (anonymous):

Rewrite with only sin x and cos x. cos 2x - sin x

OpenStudy (anonymous):

@bohotness

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

2x - sin x, how to rewrite this any helpful notes

OpenStudy (anonymous):

what??

OpenStudy (mathmate):

Use the double angle formulae: \(\rm sin(2x)=2sin(x)cos(x)\) \(\rm cos(2x)=cos^2(x)-sin^2(x)\)

OpenStudy (bohotness):

(1) First, find sin(3x) and cos(3x) in terms of sine and cosine as follows: sin(3x) = sin(2x + x) = sin(2x)cos(x) + cos(2x)sin(x), by the sine addition formula = 2sin(x)cos^2(x) + sin(x)[1 - 2sin^2(x)], by the sine/cosine double-angle formulas = 2sin(x)[1 - sin^2(x)] + sin(x)[1 - 2sin^2(x)], by the Pythagorean Identity = 2sin(x) - 2sin^3(x) + sin(x) - 2sin^3(x) = 3sin(x) - 4sin^3(x). cos(3x) = cos(2x + x) = cos(2x)cos(x) - sin(2x)sin(x), by the cosine addition formula = cos(x)[2cos^2(x) - 1] - 2sin^2(x)cos(x), by the sine/cosine double-angle formulas = cos(x)[2cos^2(x) - 1] - 2cos(x)[1 - cos^2(x)], by the Pythagorean Identity = 2cos^3(x) - cos(x) - 2cos(x) + 2cos^3(x) = 4cos^3(x) - 3cos(x). Thus: sin(3x) + cos(3x) = [3sin(x) - 4sin^3(x)] + [4cos^3(x) - 3cos(x)] = 4cos^3(x) - 4sin^3(x) + 3sin(x) - 3cos(x). (2) By the double-angle formula for cosine: cos(2x) - sin(x) = 1 - 2sin^2(x) - sin(x).

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

so the answer is cos^2x-sin^3x?

OpenStudy (bohotness):

:DDDDDDDDDDDDDDDDD

OpenStudy (mathmate):

@bohotness I don't quite get where you're going with cos(3x), etc. But the second part makes sense!

OpenStudy (anonymous):

heres my choice 1) cos^2x - sin^2x - sinx 2) cos^2x - sin^3x 3) cos^2x + sin^2x 4) cos^2x - 3sin x

jimthompson5910 (jim_thompson5910):

ybarrap answered this already http://openstudy.com/users/dglicks#/updates/54f506b0e4b09f47f4e49152

OpenStudy (anonymous):

didnt expect him to just gimme the answer like that lol thanks jim